# Finite Expression Method for Learning Dynamics on Complex Networks

### v1

@@ Abstract
Complex network data pervades various real-world domains, including physical, technological, and biological systems. Despite the prevalence of such data, predicting trends and understanding behavioral patterns in complex systems remains challenging due to poorly understood underlying mechanisms. While data-driven methods have made strides in uncovering governing equations from time series data, efforts to extract physical laws from network data are limited and often struggle with incomplete or noisy data. To address these challenges, we introduce a novel approach called the Finite Expression Method (FEX) and its fast algorithm for this learning problem on complex networks. FEX represents dynamics on complex networks using binary trees composed of finite mathematical operators. The nodes within these trees are trained through a combinatorial optimization process guided by reinforcement learning techniques. This unique configuration allows FEX to capture complex dynamics with minimal prior knowledge of the system and a small dictionary of mathematical operators. Our extensive numerical experiments demonstrate that FEX excels in accurately identifying dynamics across diverse network topologies and dynamic behaviors.
@@ 1 Introduction
Complex network data is prevalent in various real-world domains, spanning from transportation [44, 62, 72, 74] to biology [13, 10, 43] and epidemiology [49, 15, 55]. For instance, proteins frequently interact with one another to execute a wide range of functions. As with many complex systems, comprehending the underlying dynamics within these networks is crucial for analyzing and predicting their behaviors. In practice, time-series data for individual nodes within a network are often accessible. Still, extracting the dynamic interactions among these nodes to gain insights into the system’s behavior proves to be a challenging endeavor. Consequently, the physical laws governing the majority of real-world complex networks remain elusive and underexplored.
In recent years, machine learning has emerged as a promising avenue for data-driven discovery of physical laws. Notably, deep learning has demonstrated its potential to identify governing partial differential equations (PDEs) from noisy data  [50, 51, 48, 65, 39, 40, 17, 57, 3, 8, 28]. The success of these methods is attributed to the expressive power of deep learning models in capturing relationships among different variables. However, there has been limited focus on the discovery of physical laws within complex networks. Discovering dynamics on network data involves various challenges:
  • Complexity and scale: Real-world network data can be extremely large and complex, making analysis computationally expensive. In the full interaction model with $N$ nodes in the network, each node interacts with every other particle, leading to $N(N-1)$ interaction terms that need to be computed at each time, resulting a $O(N^{2})$ complexity.
@@ 1.1 Related work
1.1.1 Discoverning physical laws on network
SINDy: The Sparse Identification of Nonlinear Dynamics (SINDy) [11] algorithm aims to extract underlying governing equations from data in a interpretable and efficient way, with theoretical convergence analysis [75]. Taking time-series data as inputs, SINDy infers the underlying physical laws of the system by using Sequentially Thresholded Least Squares algorithm to choose a parsimonious model from a large library of candidate functions, usually including polynomials and trignometric functions. SINDy has been used to identify models in fluid dynamics [23], control systems [37], and neuroscience [14], etc. However, it has been shown that the vanilla SINDy performs poorly in identifying dynamics on network data.
Two-Phase SINDy: In [25], Gao and Yan proposed a two-phase procedure consisting of global regression and local fine-tuning for graph dynamics inference. This method works by building two comprehensive function libraries, $L_{F}$ and $L_{G}$, for self and interaction dynamics. Then, in the first phase, this method performs a sparse regression on the normalized input data and the libraries $L_{F}$ and $L_{G}$ to identify potential terms in the network dynamics. In the second phase, it performs topological sampling to fine-tune the coefficients of functions from the first phase. This method can identify the explicit functional forms of the system dynamics, given sufficient large function libraries. However, users often lack a priori knowledge of the candidate functions, so they assume a prior solid knowledge of the system. In addition, this method is sensitive to noisy and low-resolution data because topological samplings are used in the fine-tuning stage.
@@ 1.1.2 Fast algorithms for scientific computation
In the realm of physical sciences, it is important to develop fast and efficient algorithms to handle complex systems while keeping a high accuracy. For example, Jin et al. [36] proposed the Random Batch Methods (RBM), which splits particles in a physical system into several batches at each time to reduce the computational cost from $O(N^{2})$ to $O(N)$, where $N$ is the number of particles. In machine learning, stochastic gradient descent (SGD) [2, 9, 32] takes a random sample of the entire dataset in each iteration to compute the gradient and update model parameters. Similar to SGD, stochastic coordinate descent [46, 47] randomly samples some of the coordinates to update at each time step. In addition, randomized algorithms have been widely applied to fast matrix factorizations [18, 31, 53] and Bayesian inference [71, 52].
@@ 1.2 Our contributions
To tackle the problems seen in previous methods, we propose the finite expression method (FEX) that represents dynamics on network data by binary trees, whose nodes are mathematical operators (e.g., sin, cos, exp, etc). FEX selects the operator of each node via combinatorial optimization (CO) by a novel reinforcement learning (RL) approach. In particular, our method has major characteristics. (1) Compared to the widely used SINDy approach and its variants, FEX does not require a large library of candidate functions. In contrast, thanks to the binary tree representation, FEX can produce complicated composition of functions with a few simple mathematical operators provided by the user. (2) Our model employs a RL approach to perform operator selection in a data-driven manner (discussed in Section 3). This design has been demonstrated to promote the expressiveness of our method of identifying network dynamics. (3) To mitigate the intensive computational burden when calculating the pair-wise interaction dynamics on network data, we employ RBM [36] to accelerate the computation while keeping a reasonable accuracy. Throughout this work, we demonstrate the advantage of our method by applying it to several synthetic complex network data to extract governing equations. An overview figure of FEX to infer physical laws on network data is provided in Fig. 1.
@@ 2.1 Learning dynamics on complex networks
where $\mathbf{x}_{i}(t)\equiv\left(x_{i,1}(t),\ldots,x_{i,d}(t)\right)^{\top}$ represents node $i$’s $d$-dimensional activity and $i=1,\ldots,N$. The dynamics of $\mathbf{x}_{i}(t)$ are driven by $\mathbf{F}(\mathbf{x}_{i})\equiv\left(F_{1}\left(\mathbf{x}_{i}\right),\ldots,F_{d}\left(\mathbf{x}_{i}\right)\right)^{\top}$, denoting the self-dynamics, and $\mathbf{G}(\mathbf{x}_{i}(t),\mathbf{x}_{j}(t))\equiv\left(G_{1}\left(\mathbf{x}_{i},\mathbf{x}_{j}\right),\ldots,G_{d}\left(\mathbf{x}_{i},\mathbf{x}_{j}\right)\right)^{\top}$, representing the interaction dynamics between any pair of nodes $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$. $\mathbf{A}_{ij}$ denotes the $ij$-th entry in the $N\times N$ adjacency matrix $\mathbf{A}$, where $\mathbf{A}_{ij}=1$ if there is a connection between node $i$ and node $j$, and $\mathbf{A}_{ij}=0$ otherwise. Therefore, the dynamics are determined by self-dynamics, interaction dynamics and the topology of the network. However, in many real-world complex network systems, we do not have the information of the explicit forms of $\mathbf{F}$ and $\mathbf{G}$. Therefore, it is important to distill the functional forms of $\mathbf{F}$ and $\mathbf{G}$ from node activity data, which is the focus of our work.
However, inferring underlying dynamics from real-world time series data presents a formidable challenge, primarily due to the often unclear or only partially known governing laws. The endeavor to discover dynamics within a complex network system introduces an additional layer of complexity, attributable to the interactions among entities described by the graph topology. Moreover, numerous oscillator systems exhibit varied coupling behaviors contingent upon the coupling strength between entities and the inherent graph topology, a prime example being synchronization phenomena. In addition, the data in real-world are often sparse, noisy, and come from incomplete network topology. Hence, to demonstrate the capability of FEX to discover dynamics on complex networks, we subject FEX to a rigorous examination against three classical network dynamics: Hindmarsh-Rose (HR, $d=3$[34, 64, 70], FitzHugh-Nagumo (FHN, $d=2$) neuronal systems [19], and coupled heterogeneous Rössler oscillators ($d=3$[7], where $d$ denotes the dimension of feature for each node. Regarding to the graph topology, we demonstrate that FEX is capable of identifying governing equations on various synthetic networks, including Erdős-Rényi (ER) [27] and scale-free (SF) [1] networks. Through this extensive examination, we aim to demonstrate the effectiveness and robustness of FEX in decoding the intricate dynamics inherent in complex network systems.
@@ 2.2 FEX for learning dynamics on networks
The motivation behind FEX is to develop a robust methodology for identifying physical laws from network data. Our model depicts network dynamics using two fixed-size binary trees: one representing the self-dynamics $\mathbf{F}(\mathbf{x}_{i})$ while another one representing the interaction dynamics $\mathbf{G}(\mathbf{x}_{i},\mathbf{x}_{j})$. Each tree node denotes a simple mathematical operator—either unary (sin, cos, and exp …) or binary (+,,×+,-$\times$+ , - , ×…). The aim of the learning process is to identify the most suitable mathematical operator for each node of the tree, ensuring accurate representation of the dynamics. The input to FEX algorithm consists of time series data $\mathbf{x}_{i}$ for each node in the network, and the adjacency matrix $\mathbf{A}$ representing the network topology, as illustrated by Fig. 2.
@@ 2.3 Learning dynamics from sparse and noisy data
Spurious and missing links: In the real world, network topology may undergo evolution, making the capture of all node interactions a challenging, if not unfeasible, task. To assess the robustness of FEX against incomplete topology, we randomly either add or remove a proportion of links from the authentic network topology. The results indicate that FEX can accommodate approximately $15\%$ missing or spurious links, showcasing its robustness in handling topological inaccuracies.
Observational noise: Observational noise is another common issue in data measurement. We replicate such noise by introducing Gaussian noises to the time series data for every node within the network. The noise intensity is modulated via the signal-to-noise ratio (SNR). The findings underscore that FEX sustains a high accuracy level, even amidst a 40-dB observational noise environment. In contrast, other baseline methods such as Two-Phase SINDy [25] and ARNI [12] identified erroneous and redundant terms with low-accurate coefficients.
Through this thorough evaluation across varied scenarios and a comparative analysis with baseline models, we aim to demonstrate the robustness and stability of FEX in extracting dynamics from real-world, noisy data with potential topological inaccuracies. Further details are provided in Section 4.
@@ 3.1 Overview of FEX
FEX [42, 63] introduces a novel strategy for solving PDEs and identifying physical laws from data [35]. FEX represents the system’s dynamics using a learnable binary tree to identify dynamics on network data. This tree comprises unary operators (e.g., sin, cos, etc.) and binary operators (+,,×+,-$\times$+ , - , ×, etc.). Additionally, each unary operator incorporates learnable weights and biases, augmenting the expressiveness of the expression. Given a specific set of operators in the tree, input variables $\mathbf{x}\in\mathbb{R}^{d}$ are passed through the FEX tree via the leaves. The complete FEX expression is then derived through a preorder traversal of the tree. Unlike many symbolic regression (SR) methods that generate equations autoregressively, FEX allows users to predetermine the tree’s depth and size. This distinctive feature has been demonstrated to enhance FEX’s accuracy in solving PDEs and discovering governing equations. Subsequently, with a suitable loss function, FEX conducts CO using a RL strategy to perform operator selection, and continuous optimization to refine the parameters associated with unary operators.
@@
In particular, analyzing dynamics in complex networks brings forth several challenges, primarily the computational complexity involved. Computing particle interactions on a graph requires $O(N^{2})$ complexity, where $N$ is the number of nodes in a network. This is prohibitive for large-scale systems. We’ve employed the RBM [36] when calculating the loss function to mitigate this. Simply put, nodes in the network are grouped into random batches and particle interactions are limited to their respective batches. By integrating RBM into the CO segment of FEX training, we’ve achieved faster computational speeds, i.e., $O(N)$ computational complexity, without sacrificing the precision required to identify the accurate structure of the FEX tree.
@@ 3.3 The combinatorial optimization problem in FEX
In FEX, the loss functional $\mathcal{L}$ varies based on the problem. For this study, we adopt the least squares loss functional, as presented in [25], which is described as follows:
$\mathcal{L}(F,G)=\int_{0}^{T}\left(\left\|\text{FEX\_Tree}_{F}(\mathbf{x},t)+{\mathbf{A}}\text{FEX\_Tree}_{G}(\mathbf{x},t)-\dot{\mathbf{x}}(t)\right\|^{2}\right)\mathrm{d}t,$ (2)
[…]
$\min_{F,G\in\mathbb{S}_{\sf FEX}}\mathcal{L}(F,G),$ (3)

### v3

# A Fast Algorithm for the Finite Expression Method in Learning Dynamics on Complex Networks

@@ Abstract
Complex network data is prevalent in various real-world domains, including physical, technological, and biological systems. Despite this prevalence, predicting trends and understanding behavioral patterns in complex systems remain challenging due to poorly understood underlying mechanisms. While data-driven methods have advanced in uncovering governing equations from time series data, efforts to extract physical laws from network data are limited and often struggle with incomplete or noisy data. Additionally, they suffer from computational costs on network data, making it difficult to scale to real-world networks. To address these challenges, we introduce a novel approach called the Finite Expression Method (FEX) and its fast algorithm for learning dynamics on complex networks. FEX represents dynamics on complex networks using binary trees composed of finite mathematical operators. The nodes within these trees are trained through a combinatorial optimization process guided by reinforcement learning techniques. This unique configuration allows FEX to capture complex dynamics with minimal prior knowledge of the system and a small dictionary of mathematical operators. We also integrate a fast, stochastic algorithm into FEX, reducing the computational complexity from $O(N^{2})$ to $O(N)$. Our extensive numerical experiments demonstrate that FEX excels in accurately identifying dynamics across diverse network topologies and dynamic behaviors.
@@ 1 Introduction
Complex network data is prevalent in various real-world domains, spanning from transportation [52, 76, 87, 89] to biology [16, 9, 51] and epidemiology [57, 20, 65]. For instance, proteins frequently interact with one another to execute a wide range of functions. As with many complex systems, comprehending the underlying dynamics within these networks is crucial for analyzing and predicting their behaviors. In practice, time-series data for individual nodes within a network are often accessible. Still, extracting the dynamic interactions among these nodes to gain insights into the system’s behavior proves to be a challenging endeavor. Additionally, the interaction dynamics among particles in the network significantly increase the computational expense of identifying physical laws. Consequently, the physical laws governing the majority of real-world complex networks remain elusive and underexplored.
In recent years, machine learning has emerged as a promising avenue for data-driven discovery of physical laws. Notably, deep learning has demonstrated its potential to identify governing partial differential equations (PDEs) from noisy data  [14, 34, 58, 60, 56, 80, 45, 46, 22, 67, 3, 6, 34, 17, 18, 69, 28, 48, 84, 39, 7, 63, 71, 74, 12, 15, 68, 75, 13, 59]. The success of these methods is attributed to the expressive power of deep learning models in capturing relationships among different variables. However, deep learning methods suffer from the black-box nature of their solutions, making it difficult to interpret and explicitly identify the governing equation in mathematical form. To address this, recent works have explored symbolic regression techniques to infer governing equations, thereby obtaining an explicit mathematical representation of the system [83, 49, 79, 41, 35]. Nonetheless, there has been limited focus on the discovery of physical laws within complex networks. Discovering dynamics on network data involves various challenges:
  • Complexity and scale: Real-world network data can be extremely large and complex, making analysis computationally expensive. In the full interaction model with $N$ nodes in the network, each node interacts with every other particle, leading to N(N1)N(N-1) interaction terms that need to be computed at each time, resulting a $O(N^{2})$ complexity.
@@ 1.1 Related work
1.1.1 Discovering physical laws on network
SINDy: The Sparse Identification of Nonlinear Dynamics (SINDy) [10] algorithm aims to extract underlying governing equations from data in a interpretable and efficient way, with theoretical convergence analysis [90]. Taking time-series data as inputs, SINDy infers the underlying physical laws of the system by using Sequentially Thresholded Least Squares algorithm to choose a parsimonious model from a large library of candidate functions, usually including polynomials and trignometric functions. SINDy has been used to identify models in fluid dynamics [29], control systems [44], and neuroscience [19], etc. However, it has been shown that the vanilla SINDy performs poorly in identifying dynamics on network data. In addition, SINDy assumes direct access to the time derivative of the dynamics in the candidate function library, which is often unavailable in most real-world applications.
Two-Phase SINDy: In [31], Gao and Yan proposed a two-phase procedure consisting of global regression and local fine-tuning for graph dynamics inference. This method works by building two comprehensive function libraries, $L_{F}$ and $L_{G}$, for self and interaction dynamics. Then, in the first phase, this method performs a sparse regression on the normalized input data and the libraries $L_{F}$ and $L_{G}$ to identify potential terms in the network dynamics. In the second phase, it performs topological sampling to fine-tune the coefficients of functions from the first phase. This method can identify the explicit functional forms of the system dynamics, given sufficiently large function libraries. However, users often lack a priori knowledge of the candidate functions, so they assume a prior solid knowledge of the system. In addition, this method is sensitive to noisy and low-resolution data because topological samplings are used in the fine-tuning stage.
@@ 1.1.2 Fast algorithms for scientific computation
In the realm of physical sciences, it is important to develop fast and efficient algorithms to handle complex systems while keeping a high accuracy. For example, Jin et al. [43] proposed the Random Batch Methods (RBM), which splits particles in a physical system into several batches at each time to reduce the computational cost from $O(N^{2})$ to $O(N)$, where $N$ is the number of particles. In machine learning, stochastic gradient descent (SGD) [2, 8, 38] takes a random sample of the entire dataset in each iteration to compute the gradient and update model parameters. Similar to SGD, stochastic coordinate descent [54, 55] randomly samples some of the coordinates to update at each time step. In addition, randomized algorithms have been widely applied to fast matrix factorizations [23, 37, 62] and Bayesian inference [86, 61], etc. In spirit, our work is similar to RBM; however, instead of taking a random batch of particles for the evolution of the dynamics, we use a stochastic approach to select a random batch of particles at each step for discovering the physical laws, i.e., solving an optimization problem.
@@ 1.2 Our contributions
To tackle the problems seen in previous methods, we propose the finite expression method (FEX) that represents dynamics on network data by binary trees, whose nodes are mathematical operators (e.g., sin, cos, exp, etc). FEX selects the operator of each node via combinatorial optimization (CO) by a novel reinforcement learning (RL) approach. In particular, our method has the following advantages: (1) Compared to the widely used SINDy approach and its variants, FEX does not require a large library of candidate functions. In contrast, thanks to the binary tree representation, FEX can produce complicated composition of functions with a few simple mathematical operators provided by the user. In addition, unlike SINDy, FEX directly estimates the time derivatives from time series data without direct access to the derivatives. This is a more challenging task due to the difficulty of estimating derivatives from noisy time series data. (2) Our model employs a RL approach to perform operator selection in a data-driven manner (discussed in Section 3). This design has been demonstrated to promote the expressiveness of our method of identifying network dynamics. (3) To mitigate the intensive computational burden when calculating the pair-wise interaction dynamics on network data, we utilize a stochastic algorithm to accelerate the computation while keeping a reasonable accuracy. Throughout this work, we demonstrate the advantage of our method by applying it to several synthetic complex network data to extract governing equations. An overview figure of FEX to infer physical laws on network data is provided in Fig. 1.
@@ 2.1 Learning dynamics on complex networks
where $\mathbf{x}_{i}(t)\equiv\left(x_{i,1}(t),\ldots,x_{i,d}(t)\right)^{\top}$ represents node $i$’s $d$-dimensional feature and $i=1,\ldots,N$. The dynamics of $\mathbf{x}_{i}(t)$ are driven by $\mathbf{F}(\mathbf{x}_{i})\equiv\left(F_{1}\left(\mathbf{x}_{i}\right),\ldots,F_{d}\left(\mathbf{x}_{i}\right)\right)^{\top}$, denoting the self-dynamics, and $\mathbf{G}(\mathbf{x}_{i}(t),\mathbf{x}_{j}(t))\equiv\left(G_{1}\left(\mathbf{x}_{i},\mathbf{x}_{j}\right),\ldots,G_{d}\left(\mathbf{x}_{i},\mathbf{x}_{j}\right)\right)^{\top}$, representing the interaction dynamics between any pair of nodes $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$. $\mathbf{A}_{ij}$ denotes the $ij$-th entry in the $N\times N$ adjacency matrix $\mathbf{A}$, where $\mathbf{A}_{ij}=1$ if there is a connection between node $i$ and node $j$, and $\mathbf{A}_{ij}=0$ otherwise. Therefore, the dynamics are determined by self-dynamics, interaction dynamics and the topology of the network. However, in many real-world complex network systems, we do not have the information of the explicit forms of $\mathbf{F}$ and $\mathbf{G}$. Therefore, it is important to distill the functional forms of $\mathbf{F}$ and $\mathbf{G}$ from node activity data, which is the focus of our work.
However, inferring underlying dynamics from real-world time series data presents a formidable challenge, primarily due to the often unclear or only partially known governing laws. The endeavor to discover dynamics within a complex network system introduces an additional layer of complexity, attributable to the interactions among entities described by the graph topology. Moreover, numerous oscillator systems exhibit varied coupling behaviors contingent upon the coupling strength between entities and the inherent graph topology, a prime example being synchronization phenomena. In addition, the data in real-world are often sparse, noisy, and come from incomplete network topology. Hence, to demonstrate the capability of FEX to discover dynamics on complex networks, we subject FEX to a rigorous examination against three classical network dynamics: Hindmarsh-Rose (HR, d=3d=3[40, 78, 85], FitzHugh-Nagumo (FHN, $d=2$) neuronal systems [24], and coupled heterogeneous Rössler oscillators (d=3d=3[5], where $d$ denotes the dimension of feature for each node. Regarding to the graph topology, we demonstrate that FEX is capable of identifying governing equations on various synthetic networks, including Erdős-Rényi (ER) [33] and scale-free (SF) [1] networks. Through this extensive examination, we aim to demonstrate the effectiveness and robustness of FEX in decoding the intricate dynamics inherent in complex network systems. Details are elaborated more in Section 4.
@@ 2.2 FEX for learning dynamics on networks
The motivation behind FEX is to develop a robust methodology for identifying physical laws from network data. Our model depicts network dynamics using two fixed-size binary trees: one representing the self-dynamics $\mathbf{F}(\mathbf{x}_{i})$ while another one representing the interaction dynamics $\mathbf{G}(\mathbf{x}_{i},\mathbf{x}_{j})$. Each tree node denotes a simple mathematical operator—either unary (sin, cos, and exp …) or binary ($+,-,\times$…). The aim of the learning process is to identify the most suitable mathematical operator for each node of the tree, ensuring accurate representation of the dynamics. The input to FEX algorithm consists of time series data $\mathbf{x}_{i}$ for each node in the network, and the adjacency matrix $\mathbf{A}$ representing the network topology, as illustrated by Fig. 2.
@@ 2.3 Learning dynamics from sparse and noisy data
Spurious and missing links: In real-world applications, the evolving network topology makes capturing all node interactions a challenging task. To assess the robustness of FEX against incomplete topology, we randomly either add or remove a proportion of links from the authentic network topology. The results indicate that FEX can accommodate approximately $15\%$ missing or spurious links, showcasing its robustness in handling topological inaccuracies.
Observational noise: Observational noise is another common issue in data measurement. We replicate such noise by introducing Gaussian noises to the time series data for every node within the network. The noise intensity is modulated via the signal-to-noise ratio (SNR). The findings underscore that FEX sustains a high accuracy level, even amidst a 40-dB observational noise environment. In contrast, other baseline methods such as Two-Phase SINDy [31] and ARNI [11] identified erroneous and redundant terms with low-accuracy coefficients.
Through this thorough evaluation across varied scenarios and a comparative analysis with baseline models, we aim to demonstrate the robustness and stability of FEX in extracting dynamics from real-world, noisy data with potential topological inaccuracies. Further details are provided in Supplementary Information.
@@ 3.1 Overview of FEX
FEX [50, 77] introduces a novel strategy for solving PDEs and identifying physical laws from data [42]. FEX represents the system’s dynamics using a learnable binary tree to identify dynamics on network data. This tree comprises unary operators (e.g., sin, cos, etc.) and binary operators ($+,-,\times$, etc.). Additionally, each unary operator incorporates learnable weights and biases, augmenting the expressiveness of the expression. Given a specific set of operators in the tree, input variables $\mathbf{x}\in\mathbb{R}^{d}$ are passed through the FEX tree via the leaves. The complete FEX expression is then derived through a preorder traversal of the tree. Unlike many symbolic regression (SR) methods that generate equations autoregressively, FEX allows users to predetermine the tree’s depth and size. This distinctive feature has been demonstrated to enhance FEX’s accuracy in solving PDEs and discovering governing equations. Subsequently, with a suitable loss function, FEX conducts CO using a RL strategy to perform operator selection, and continuous optimization to refine the parameters associated with unary operators.
@@
In particular, analyzing dynamics in complex networks brings forth several challenges, primarily the computational complexity involved. Computing particle interactions on a graph requires $O(N^{2})$ complexity, where $N$ is the number of nodes in a network. This is prohibitive for large-scale systems. We utilize a stochastic algorithm when calculating the loss function to mitigate this. Simply put, nodes in the network are grouped into random batches and particle interactions are limited to their respective batches. With this strategy, we achieve faster computational speeds, i.e., $O(N)$ computational complexity, without sacrificing the precision required to identify the accurate structure of the FEX tree.
@@ 3.3 The combinatorial optimization problem in FEX
In FEX, the loss functional $\mathcal{L}$ varies based on the problem. For this study, we adopt the least squares loss functional on each dimension of the dynamics, as presented in [31], which is described as follows:
$\mathcal{L}(\theta_{F},\theta_{G})=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left\|\frac{d\mathbf{x}_{i}(t)}{dt}-\left(\mathbf{F}(\mathbf{x}_{i}(t);\theta_{F})+\sum_{j=1}^{N}\mathbf{A}_{ij}\mathbf{G}(\mathbf{x}_{i}(t),\mathbf{x}_{j}(t);\theta_{G})\right)\right\|_{2}^{2},$ (2)
where $\mathbf{A}$ is the adjacency matrix of the network, $\mathbf{F}(\mathbf{x}_{i}(t);\theta_{F})$ denotes the self-dynamics modeled by a FEX tree parameterized by $\theta_{F}$, $\mathbf{G}(\mathbf{x}_{i}(t),\mathbf{x}_{j}(t);\theta_{G})$ denotes the interaction dynamics by another FEX tree parameterized by $\theta_{G}$. In FEX, the solution is found by solving the combinatorial optimization problem
$\min_{\mathbf{F},\mathbf{G}\in\mathbb{S}_{\sf FEX}}\mathcal{L}(\mathbf{F},\mathbf{G}),$ (3)